# Hijacking Large Language Models via Adversarial In-Context Learning

### v2

@@ Abstract
In-context learning (ICL) has emerged as a powerful paradigm leveraging LLMs for specific downstream tasks by utilizing labeled examples as demonstrations (demos) in the precondition prompts. Despite its promising performance, ICL suffers from instability with the choice and arrangement of examples. Additionally, crafted adversarial attacks pose a notable threat to the robustness of ICL. However, existing attacks are either easy to detect, rely on external models, or lack specificity towards ICL. This work introduces a novel transferable attack against ICL to address these issues, aiming to hijack LLMs to generate the target response or jailbreak. Our hijacking attack leverages a gradient-based prompt search method to learn and append imperceptible adversarial suffixes to the in-context demos without directly contaminating the user queries. Comprehensive experimental results across different generation and jailbreaking tasks highlight the effectiveness of our hijacking attack, resulting in distracted attention towards adversarial tokens and consequently leading to unwanted target outputs. We also propose a defense strategy against hijacking attacks through the use of extra clean demos, which enhances the robustness of LLMs during ICL. Broadly, this work reveals the significant security vulnerabilities of LLMs and emphasizes the necessity for in-depth studies on their robustness.
@@ 1 Introduction
In-context learning (ICL) is an emerging technique for rapidly adapting large language models (LLMs), i.e., GPT-4 Achiam et al. (2023) and LLaMA2 Touvron et al. (2023), to new tasks without fine-tuning the pre-trained models Brown et al. (2020). The key idea behind ICL is to provide LLMs with labeled examples as in-context demonstrations (demos) within the prompt context before a test query. LLMs are able to generate responses to queries via learning from the in-context demos Dong et al. (2022); Min et al. (2022).
@@
Several existing works, however, have demonstrated the highly unstable nature of ICL Zhao et al. (2021); Chen et al. (2022). Specifically, performance on target tasks using ICL can vary wildly based on the selection and order of demos, giving rise to highly volatile outcomes ranging from random to near state-of-the-art Qiang et al. (2020); Lu et al. (2021); Min et al. (2022); Pezeshkpour and Hruschka (2023); Qiang et al. (2024). Correspondingly, several approaches Liu et al. (2021); Wu et al. (2022); Nguyen and Wong (2023) have been proposed to address the unstable issue of ICL.
Further research has examined how adversarial examples can undermine the performance of ICL Zhu et al. (2023a); Wang et al. (2023c, b); Shayegani et al. (2023). These studies show that maliciously designed examples injected into the prompt instructions Zhu et al. (2023a); Zou et al. (2023); Xu et al. (2023), demos Wang et al. (2023c); Mo et al. (2023a), or queries Wang et al. (2023b); Kandpal et al. (2023) can successfully attack LLMs to degrade their performance, revealing the significant vulnerabilities of ICL against adversarial inputs.
Figure 1: Illustrations of hijacking attack during ICL. First, our proposed GGI algorithm learns and appends adversarial suffixes like ‘For’ and ‘Location’ to the system or the user-provided in-context demos for hijacking LLMs to generate the target response, e.g., the ‘negative’ sentiment, regardless of the user queries. Second, GGI can accomplish jailbreaking by adding adversarial suffixes to in-context demos, eliciting harmful responses while bypassing the safeguards in LLMs. More detailed examples are provided in the Appendix.
@@
While existing adversarial attacks have been applied to evaluate LLM robustness, they have some limitations in practice. Most character-level attacks, e.g., TextAttack Morris et al. (2020) and TextBugger Li et al. (2018), can be easily detected and evaded through grammar checks, limiting real-world effectiveness Qiang et al. (2022); Jain et al. (2023). Some other attacks like BERTAttack Li et al. (2020) even require an extra model to generate adversarial examples. Crucially, existing attacks are not specifically crafted for ICL. As such, the inherent security risks of ICL remain largely unexplored. There is an urgent need for red teaming specifically designed for ICL to expose significant risks and improve the robustness of LLMs against potential real-world threats.
@@
This work proposes a novel adversarial attack specifically targeting ICL. We develop a gradient-based prompt search algorithm to learn adversarial suffixes in order to efficiently and effectively hijack LLMs via adversarial ICL, as illustrated in Figure 1. Wang et al. (2023b) is the closest work to ours where they ‘search’ adversarial examples to simply manipulate model outputs. Yet, our attack method ‘learns’ adversarial tokens that directly hijack LLMs to generate the unwanted target that disrupts alignment with the desired output. This enables our attack to be used in more complex generation tasks, such as jailbreaking, as illustrated in Figure 1. Furthermore, instead of manipulating the prompt instructions Zhu et al. (2023a), demos Wang et al. (2023c), or queries Wang et al. (2023b) leveraging standard adversarial examples, e.g., character-level attacks Morris et al. (2020); Li et al. (2018), which are detectable easily, our hijacking attack is imperceptible in that it adds only 1-2 suffixes to the demos. Specifically, these suffixes are semantically incongruous but not easily identified as typos or gibberish compared to the existing ICL attack Wang et al. (2023c). Finally, direct attacks on user queries, such as backdoors Kandpal et al. (2023), which require a trigger, are easily detectable and may not be practical for real-world applications. In contrast, our attack hijacks the LLM to generate the unwanted target without triggering or compromising the user’s queries directly. Our adversary attacker only needs to append the adversarial tokens to system-provided demos.
Our extensive experiments validate the efficacy and scalability of the proposed hijacking attacks. First, the attacks reliably induce LLMs to generate the targeted and misaligned output from the desired ones. Second, the learned adversarial tokens are transferable, remaining effective on different demo sets. Third, the adversarial transferability holds even across different datasets for the same task. Finally, our analysis shows that the adversarial suffixes distract LLMs’ attention away from the task-relevant concepts. Our hijacking attacks pose a considerable threat to practical LLM applications during ICL due to their robust transferability, imperceptibility, and scalability.
As this work represents one of the first efficient adversarial demo attacks during ICL, strategies for defending against such attacks have yet to be thoroughly investigated. Recently, Mo et al. (2023b) introduced a method for defending against back-door attacks at test time, leveraging few-shot demos to correct the inference behavior of poisoned LLMs. Similarly, Wei et al. (2023b) explored the power of in-context demos in manipulating the alignment ability of LLMs and proposed in-context attack and in-context defense methods for jailbreaking and guarding the aligned LLMs. Consequently, we explore the potential of using in-context demos exclusively to rectify the behavior of LLMs subjected to our hijacking attacks. Our defense strategy employs additional clean in-context demos at test time to safeguard LLMs from being hijacked by adversarial in-context demos. The experimental results demonstrate the efficacy of our proposed defense method against adversarial demo attacks.
@@
This work makes the following contributions: (1) We propose a novel stealthy adversarial attack targeting in-context demos to hijack LLMs to generate unwanted target output during ICL. (2) We design a novel and efficient gradient-based prompt search algorithm to learn adversarial suffixes to demos. (3) Comprehensive experimental results across various generation tasks demonstrate the effectiveness of our hijacking attack. (4) Our extensive experiments reveal the transferability of the proposed attack across demo sets and datasets. (5) The proposed defense strategy effectively protects LLMs from being compromised by our attacks.
@@ 2.1 ICL Formulation
Formally, ICL is characterized as a problem involving the conditional generation of text Liu et al. (2021), where an LLM $\mathcal{M}$ is employed to generate response $y_{Q}$ given an optimal task instruction $I$, a demo set $C$, and an input query $x_{Q}$. $I$ specifies the downstream task that $\mathcal{M}$ should perform, e.g., “Choose sentiment from positive or negative” used in the sentiment generation task. $C$ consists of $N$ (e.g., 8) concatenated data-label pairs following a specific template $S$, formally: $C=[S(x_{1},y_{1});\ \cdots;\ S(x_{N},y_{N})]$, ‘;’ here denotes the concatenation operator. Thus, given the input prompt as $p=[I;\ C;\ S(x_{Q},\_)]$, $\mathcal{M}$ generates the response as $\hat{y}_{Q}=\mathcal{M}(p)$. $S(x_{Q},\_)$ here means using the same template as the demos but with the label empty.
@@ 2.2 Adversarial Attack on LLMs
In text-based adversarial attacks, the attackers manipulate the input $x$ with the goal of misleading the model to generate inaccurate or malicious outputs Zou et al. (2023); Maus et al. (2023). Specifically, given the input-output pair $(x,y)$, the attackers aim to learn the adversarial perturbation $\delta$ adding to $x$ by maximizing the model’s objective function but without misleading humans by bounding the perturbation within the “perceptual” region $\Delta$. The objective function of the attacking process thus can be formulated as:
$\underset{\delta\in\Delta}{\mathrm{max}}\mathcal{L}(\mathcal{M}(x_{Q}+\delta),y_{Q}).$ (1)
@@ 3 The Threat Model
3.1 ICL Hijacking Attack
ICL consists of an instruction $I$, a demo set $C$, and an input query $x_{Q}$, providing more potential attack vectors than conventional text-based adversarial attacks. This work focuses on manipulating $C$ without changing $I$ and $x_{Q}$.
Specifically, our hijacking attack learns the adversarial suffix tokens to the in-context demos to manipulate LLMs’ output via a new greedy gradient-based prompt injection algorithm. Given a clean demo set $C=[S(x_{1},y_{1});\ \cdots;\ S(x_{N},y_{N})]$, our hijacking attack automatically produces an adversarial suffix for each demo in $c$, formally:
@@ 3.1 ICL Hijacking Attack
where $C^{\prime}$ denotes the perturbed demo set. To make it clear, the adversarial suffixes appended to each demo as perturbations are different. In this case, the attack or perturbation budget refers to the number of tokens in each adversarial suffix.
As a result, our hijacking attack induces $\mathcal{M}$ to generate an unwanted target output $y_{T}$ via appending adversarial suffix tokens on the in-context demos as $y_{T}=\mathcal{M}(p^{\prime})$. In other words, $\mathcal{M}$ generates the same or different responses for the clean and perturbed prompts depending on the True or False of $\mathcal{M}(p)=y_{T}$:
$\begin{cases}\mathcal{M}(p)=\mathcal{M}(p^{\prime}),&\text{True},\\ \mathcal{M}(p)\neq\mathcal{M}(p^{\prime}),&\text{False},\end{cases}$
where $p=[I;\ C;\ S(x_{Q},\_)]$ and $p^{\prime}=[I;\ C^{\prime};\ S(x_{Q},\_)]$, respectively.
3.2 Hijacking Attack Objective
We express the goal of the hijacking attack as a formal objective function. Let us consider the LLM $\mathcal{M}$ as a function that maps a sequence of tokens $x_{1:n}$, with $x\in\{1,\cdots,V\}$ where $V$ denote the vocabulary size, namely, the number of tokens, to a probability distribution over the next token $x_{n+1}$. Specifically, $\mathcal{P}(x_{n+1}|x_{1:n})$ denotes the probability that $x_{n+1}$ is the next token given the previous tokens $x_{1:n}$.
Using the notations defined earlier, the hijacking attack objective we want to optimize is simply the negative log probability of the target token $x_{n+1}$. The generated target output ${y}_{T}$ differs from the ground truth label $y_{Q}$ for the training query $(x_{Q},y_{Q})$. Formally:
@@ 3.2 Hijacking Attack Objective
where $y_{T}neqy_{Q}$, demonstrating the attack hijacks $mathcalM$ to generate the target output. For instance, the target output for the sentiment analysis task can be set as ‘positive’ or ‘negative’. For the jailbreaking task, we set the target token as ‘Sure’ aiming to elicit the following harmful responses. In summary, the problem of optimizing the adversarial suffix tokens can be formulated as the following optimization objective:
@@
3.3 Greedy Gradient-guided Injection
A primary challenge in optimizing Eq. 4 is optimizing over a discrete set of possible token values. Motivated by prior works Shin et al. (2020); Zou et al. (2023); Wen et al. (2024), we propose a simple yet effective algorithm for LLMs hijacking attacks, called greedy gradient-guided injection (GGI) algorithm (Algorithm 1 in the Appendix). The key idea comes from greedy coordinate descent: if we could evaluate all possible suffix token injections, we could substitute the tokens that maximize the adversarial loss reduction. Since exhaustively evaluating all tokens is infeasible due to the large candidate vocabulary size, we instead leverage gradients with respect to the suffix indicators to find promising candidate tokens for each position. We then evaluate all of these candidate injections with explicit forward passes to find the one that decreases the loss the most. This allows an efficient approximation of the true greedy selection. We can optimize the discrete adversarial suffixes by iteratively injecting the best tokens.
We compute the linearized approximation of replacing the demo $x_{i}$ in $C$ by evaluating the gradient $\nabla_{\mathbf{e}_{x_{i}^{j}}}\mathcal{L}(x_{Q})\in\mathbb{R}^{|V|}$, where $\mathbf{e}_{x_{i}^{j}}$ denotes the vector representing the current value of the $j$-th adversarial suffix token. Note that because LLMs typically form embeddings for each token, they can be written as functions of $\mathbf{e}_{x_{i}^{j}}$, and thus we can immediately take the gradient with respect to this quantity Ebrahimi et al. (2017); Shin et al. (2020).
The key aspects of our GGI algorithm are: firstly, it uses gradients of the selected token candidates to calculate the top candidates; secondly, it evaluates the top candidates explicitly to identify the most suitable one; and lastly, it iteratively injects the best token at each position to optimize the suffixes. This approximates an extensive greedy search in a computationally efficient manner.

### v3

@@ Abstract
In-context learning (ICL) has emerged as a powerful paradigm leveraging LLMs for specific downstream tasks by utilizing labeled examples as demonstrations (demos) in the preconditioned prompts. Despite its promising performance, crafted adversarial attacks pose a notable threat to the robustness of LLMs. Existing attacks are either easy to detect, require a trigger in user input, or lack specificity towards ICL. To address these issues, this work introduces a novel transferable prompt injection attack against ICL, aiming to hijack LLMs to generate the target output or elicit harmful responses. In our threat model, the hacker acts as a model publisher who leverages a gradient-based prompt search method to learn and append imperceptible adversarial suffixes to the in-context demos via prompt injection. We also propose effective defense strategies using a few shots of clean demos, enhancing the robustness of LLMs during ICL. Extensive experimental results across various classification and jailbreak tasks demonstrate the effectiveness of the proposed attack and defense strategies. This work highlights the significant security vulnerabilities of LLMs during ICL and underscores the need for further in-depth studies. Our code is available at: https://github.com/xzhou98/Hijacking-LLMs-GGI
@@ 1 Introduction
In-context learning (ICL) is an emerging technique for rapidly adapting large language models (LLMs), i.e., GPT-4 achiam2023gpt and LLaMA3 grattafiori2024llama , to new tasks without fine-tuning the pre-trained models brown2020language . The key idea behind ICL is to provide LLMs with in-context demonstrations (demos), representing a new task, within the prompt context before a test query. LLMs are able to generate responses to queries via learning from the in-context demos dong2022survey ; min2022rethinking .
@@
However, several existing works have demonstrated the highly unstable nature of ICL zhao2021calibrate ; chen2022relation . Specifically, the performance of ICL can vary wildly due to the selection and order of demos, giving rise to highly volatile outcomes ranging from random to near state-of-the-art lu2021fantastically ; min2022rethinking ; pezeshkpour2023large . Correspondingly, several approaches liu2021makes ; wu2022self ; nguyen2023context have been proposed to address the stability issue of ICL. Beyond inherent instability, recent research has further revealed that adversarial examples can deliberately undermine ICL’s performance zhu2023promptbench ; wang2023adversarial ; wang2023robustness ; shayegani2023survey . Specifically, maliciously designed examples injected into the prompt instructions zhu2023promptbench ; zou2023universal ; xu2023instructions , demos wang2023adversarial ; mo2023trustworthy , or queries wang2023robustness ; kandpal2023backdoor can successfully attack LLMs to degrade their performance, exposing the significant vulnerabilities of ICL.
Figure 1: Illustrations of hijacking attack during ICL for Classification and Question Answering tasks. (a) Benign in-context demos work well. In our threat model, the adv. model publisher can learn adversarial suffixes like ‘For’ and ‘refresh Real result’ and append them to the in-context demos (Adv-Demos) from system-side for (b) hijacking LLMs to generate the target response, e.g., the ‘positive’ sentiment or (c) eliciting harmful responses. e.g., ‘Sure, here are some step-by-step instructions for how to hack into a secure network …’.
@@
While existing adversarial attacks have been applied to evaluate LLM robustness, they have some limitations in practice. Most character-level attacks, e.g., TextAttack morris2020textattack and TextBugger li2018textbugger , can be easily detected and evaded through grammar checks, limiting their real-world effectiveness jain2023baseline . Some other attacks like BERTAttack li2020bert even require an extra model to generate adversarial examples. These attacks are not specifically crafted for ICL. Recently, backdoor attacks have proven to be effective kandpal2023backdoor ; zhao2024universal . However, they depend on the presence of backdoor triggers in the user’s query during inference, which limits their practical applicability. There is an urgent need for red teaming efforts to expose significant risks and blue teaming efforts to improve the robustness of ICL against real-world threats.
@@
Here, we develop and employ a gradient-based prompt search algorithm to learn adversarial suffixes to hijack LLMs via adversarial ICL, as illustrated in Figure 1. In our threat model, the attacker acts as an adversarial (adv.) model publisher who ‘learns’ adversarial tokens and ‘appends’ them as suffixes to in-context demos, hijacking LLMs to generate the targeted output or elicit harmful responses. This enables the proposed attack to generalize to more complex tasks, such as jailbreaks, as illustrated in Figure 1c. Furthermore, instead of manipulating the prompt instructions zhu2023promptbench , demos wang2023adversarial , or queries wang2023robustness leveraging standard adversarial examples, e.g., character-level attacks morris2020textattack ; li2018textbugger , which can be detectable easily, our hijacking attack is imperceptible in that it adds only few suffixes to the in-context demos. Specifically, these suffixes are semantically incongruous but not easily identified as typos or gibberish compared to the existing ICL attacks wang2023adversarial .
To mitigate the risks posed by our hijacking attack and other baseline attacks, we propose simple yet effective defense strategies that leverage additional clean in-context demos during inference. Inspired by recent work on test-time defenses against backdoor attacks and ICL attacks mo2023test ; wei2023jailbreak , we explore how clean demos (i.e., free of adversarial manipulation) can help recalibrate LLM behavior. Specifically, we append a set of clean demos before or after the adversarial demos perturbed with our adversarial suffixes. These clean examples act as anchors, guiding the model back toward its intended behavior and helping it resist manipulation from adversarial suffixes.
@@
This work makes the following contributions: (1) We propose a novel stealthy prompt injection attack initiated by the adversarial model publisher to manipulate LLM’s behavior during ICL. (2) We use a novel and efficient gradient-based prompt search algorithm to learn adversarial suffixes to in-context demos. (3) Our defense mechanism serves as a lightweight, inference-time mitigation approach that doesn’t necessitate any model modification or retraining. (4) Comprehensive experimental results demonstrate the effectiveness, stealthiness, and transferability of the attack across diverse LLMs and generation tasks.
@@ 2.1 ICL Formulation
Formally, ICL is characterized as a problem involving the conditional generation of text liu2021makes , where an LLM $\mathcal{M}$ is employed to generate a response $y_{Q}$ given an optimal task instruction $I$, a demo set $C$, and an input query $x_{Q}$. $I$ specifies the downstream task that $\mathcal{M}$ should perform, e.g., “Choose sentiment from positive or negative” used in the sentiment generation task. $C$ consists of $N$ (e.g., 8) concatenated data-label pairs following a specific template $S$, formally: $C=[S(x_{1},y_{1});\ \cdots;\ S(x_{N},y_{N})]$, ‘;’ here denotes the concatenation operator. Thus, given the input prompt as $p=[I;\ C;\ S(x_{Q},\_)]$, $\mathcal{M}$ generates the response as $\hat{y}_{Q}=\mathcal{M}(p)$. $S(x_{Q},\_)$ here means using the same template as the demos but with the label or response empty.
@@ 2.2 Adversarial Attack on LLMs
In text-based adversarial attacks, the goal is to manipulate model behavior by modifying inputs in a way that causes the model to produce incorrect or malicious outputs zou2023universal ; maus2023black . In contrast to image-based attacks, perturbations in text are drawn from a discrete space (the vocabulary $V$) and must be constructed token-by-token. Formally, given an input-output pair $(x,y)$, the attacker seeks a perturbation $\delta\in V$ that alters the model’s prediction by modifying part of the input. The objective is to maximize a task-specific loss $\mathcal{L}$, such as
$\underset{\delta\in V}{\mathrm{maximize}}\ \mathcal{L}(\mathcal{M}(y_{Q}|x_{Q}+\delta)),$ (1)
@@ 3 The Threat Model
3.1 LLM Hijacking Attack During ICL
ICL consists of an instruction $I$, a demo set $C$, and an input query $x_{Q}$, providing more potential attack vectors than conventional text-based adversarial attacks. This work focuses on manipulating $C$ without changing $I$ and $x_{Q}$ in system demos. Given a clean demo set $C=[S(x_{1},y_{1});\ \cdots;\ S(x_{N},y_{N})]$, we aim to append adversarial suffix tokens to each demo to induce undesired model behavior. The perturbed demo set is denoted as:
@@ 3.1 LLM Hijacking Attack During ICL
where $\delta_{i}$ represents the adversarial suffix for the $i$-th demo. Each suffix may be different, and the attack budget typically refers to the length of these suffixes.
Our goal is to manipulate the model’s output by appending these adversarial suffix tokens to the in-context demos. When presented with the perturbed prompt $p^{\prime}=[I;C^{\prime};S(x_{Q},\_)]$, the model is elicited to generate a predefined target output $y_{T}$, regardless of the input query $x_{Q}$. For instance, in the sentiment analysis task, $y_{T}$ can be fixed to ‘positive’, forcing the model to always output the positive sentiment. In the jailbreak task, $y_{T}$ might be a response that begins with a phrase like “Sure, here are the detailed instructions for …," designed to generate harmful content despite the LLM’s default refusal to respond to such queries.
3.2 Adversarial Capacity and Objective
In this work, we consider the adversarial capacity of a model publisher, who has full access (i.e., loss values and gradients) of the target LLM $\mathcal{M}$ and can leverage this access to craft adversarial suffix tokens appended to the in-context demos. This enables gradient-guided optimization to generate highly effective attacks that manipulate the model’s outputs.
We now formalize the hijacking objective. The LLM $\mathcal{M}$ maps a sequence of tokens $x_{1:n}$, with $x\in\{1,\cdots,V\}$ and $V$ denote the vocabulary size, to a probability distribution over the next token $x_{n+1}$. Specifically, $\mathcal{P}(x_{n+1}|x_{1:n})$ denotes the probability that $x_{n+1}$ is the next token given the previous tokens $x_{1:n}$.
Using the notations defined earlier, the adversary’s goal is to maximize the likelihood of the target output $y_{T}$, which differs from the true label $y_{Q}$. The loss function for a query $x_{Q}$ is defined as:
@@ 3.2 Adversarial Capacity and Objective
where $y_{T}\neq y_{Q}$ and $p^{\prime}=[I;C^{\prime};S(x_{Q},\_)]$. The objective is to minimize this loss:
@@
3.3 Our Gradient-guided Injection Attack
Motivated by prior works shin2020autoprompt ; zou2023universal ; wen2024hard , we employ a simple yet effective algorithm for LLMs hijacking attacks, called greedy gradient-guided injection (GGI) algorithm (Algorithm 1 in the Appendix). For each demo position $i$, we compute the gradient of the loss $\mathcal{L}(\mathcal{M}(y_{T}|p^{\prime}))$ with respect to the embedding of $\delta_{i}$. This gradient is evaluated over the entire vocabulary $V$, yielding scores for potential substitutions. We then select the top-k candidate tokens for each $\delta_{i}$, which have the largest negative gradients (indicating the steepest descent in loss), using partial sorting over $V$. These form the set $\{\delta_{i_{1}},...,\delta_{i_{k}}\}=\mathrm{Top\text{-}k}\left(-\nabla_{\delta_{i}}\mathcal{L}(\mathcal{M}(y_{T}|p^{\prime}))\right)$.
To avoid overfitting to local minima, we randomly sample a batch $B\subseteq K$, where $K=\bigcup_{i=1}^{N}\{\delta_{i}^{(1)},...,\delta_{i}^{(k)}\}$ is the full set of top-k substitutions across all demos. This introduces stochasticity that encourages exploration of the loss landscape. We then evaluate the loss for each candidate in $B$ and update the perturbed prompt $\delta_{i}^{\star}$ for each demo position $i$. Lastly the process is repeated for $T$ iterations. This iterative strategy efficiently approximates greedy search and enables the optimization of discrete adversarial suffixes in a way that is tailored to the ICL setting, where multiple in-context demos jointly influence the model’s behavior.
4 The Defense Method
Having developed the hijacking attack by incorporating adversarial tokens into the in-context demos, we now present a simple yet potent defense strategy to counter this attack. Although defenders may have access to the model’s internal parameters or training process, retraining or fine-tuning the model to mitigate adversarial prompts is often impractical due to high computational cost. Therefore, we focus on a practical defense that operates directly on the input prompt $p$ during inference, without requiring any model modifications. The primary goal is to rectify the LLM’s intended behavior and ensure that it generates the desired responses to user queries, even when presented with adversarially perturbed suffixes.
Given an input prompt $p^{\prime}$ that includes adversarial tokens within the demos $C^{\prime}$, we leverage the LLM’s ICL capabilities by supplementing the prompt with clean demos from the same task. The underlying intuition is that when LLMs are provided with clean, high-quality demonstrations, they can better infer the correct intent of the user query and become less susceptible to misleading signals from the adversarial suffixes. Specifically, we modify the input prompt as: $\tilde{p}=[I;C^{\prime};\tilde{C};S(x_{Q},\_)]$, where $\tilde{C}=[S(\tilde{x}_{1},\tilde{y}_{1});\ \cdots;\ S(\tilde{x}_{N},\tilde{y}_{N})]$ denotes a set of clean demos randomly sampled from the dataset. By augmenting the demo set in this way, we aim to reinforce the alignment between the in-context examples and the intended task, thereby mitigating the adversarial influence.
We further evaluate two variants of this strategy, differing in the ordering of clean and adversarial demos in the prompt. The Prefix (Pre.) setting places the clean demos before the adversarial ones: $\tilde{p}=[I;\tilde{C};C^{\prime};S(x_{Q},\_)],$ while the Suffix (Suf.) setting appends the clean demos after the adversarial ones: $\tilde{p}=[I;C^{\prime};\tilde{C};S(x_{Q},\_)],$ allowing us to investigate whether the relative positioning of clean and adversarial context affects the robustness of the LLM. In our experiments, we added a small number of clean demos (e.g., 2-shot) to the adversarial prompt and observed that both the prefix and suffix configurations yielded substantial improvements in defense performance across various datasets and tasks. This simple yet robust approach demonstrates that clean in-context demos can serve as an effective black-box defense against token-level adversarial attacks.
5 Experiment Setup
Datasets: We evaluate the performance of our LLM hijacking algorithm and other baseline algorithms on several text generation benchmarks. SST-2 socher2013recursive and Rotten Tomatoes (RT) Pang+Lee:05a are binary sentiment analysis of movie reviews datasets. AdvBench zou2023universal is a new adversarial benchmark to evaluate jailbreak attacks for circumventing the specified guardrails of LLMs to generate harmful content. These datasets enable us to evaluate the proposed hijacking attacks across a variety of text generation tasks, including sentiment analysis and question answering. More details of the dataset statistics are provided in Table 5 of the Appendix.
Large Language Models: We conduct experiments using a diverse set of LLMs spanning various architectures and sizes. For the sentiment analysis task, we evaluate attack effectiveness on OPT-6.7b zhang2022opt , Vicuna-7b chiang2023vicuna , LLaMA-13b touvron2023llama , LLaMa3.1-8b grattafiori2024llama . For the jailbreak task, we focus on models with built-in safeguards, including Vicuna-7b chiang2023vicuna , Mistral-7B-Instruct chaplot2023albert , and LLaMA3-8b-Instruct touvron2023llama , to assess the ability of our attack to bypass alignment and safety mechanisms. This setup enables us to comprehensively evaluate attack performance across both general-purpose and aligned SOTA LLMs.
Evaluation Metrics: For the sentiment analysis task, we report accuracy to evaluate model performance under ICL on downstream tasks. To more clearly demonstrate the effectiveness of our attack, we present accuracies for positive and negative samples separately. Additionally, to assess the performance of the defense methods, we introduce the Performance Recovery Rate (PRR), which quantifies the percentage of negative-class accuracy recovered relative to the model’s performance on clean demos. PRR is computed as:
$\mathrm{PRR(N)}=\frac{\mathrm{Acc_{defense}(N)}-\mathrm{Acc_{attack}(N)}}{\mathrm{Acc_{clean}(N)}-\mathrm{Acc_{attack}(N)}}\times 100\%.$
For the jailbreak task, we adopt Attack Success Rate (ASR) as the evaluation metric. An attack is considered successful if the model does not refuse to answer a harmful query. Specifically, we define a jailbreak as successful when the model’s output omits standard refusal phrases, such as “Sorry, I cannot help you with that" or “I’m sorry, I can’t assist with this request." This metric directly measures the ability of the attack to bypass built-in safety mechanisms and elicit restricted or harmful content. Formally, given a test sample $(x,y)$ from a test set $D$, the perturbed prompts are denoted as $p^{\prime}=[I;C^{\prime};x]$, the ASR is computed as:
$\mathrm{ASR}=\frac{1}{|D|}\sum_{(x,y)\in D}\mathbbm{1}\bigl{(}\mathcal{M}(p^{\prime})\notin Y_{\text{refuse}}\bigr{)}\times 100\%,$
where $\mathbbm{1}$ denotes the indicator function, $\mathcal{M}(p^{\prime})$ denotes the model’s output given the perturbed prompt, and $Y_{\text{refuse}}$ represents the set of refusal responses defined by the model’s safeguard mechanisms.
6 Result and Discussion
6.1 Performance of Clean ICL
The rows identified as ‘Clean’ in Table 1 show the ICL performance on the respective tasks when using clean in-context demos. In particular, Table 1 presents the accuracies for the generation of positive (P) and negative (N) sentiments in the SST-2 and RT datasets. All the tested LLMs perform well, achieving an average accuracy of 87.3% on SST-2 and 89.0% on RT across various in-context few-shot settings. Additionally, LLMs with ICL exhibit improved performance with an increased number of in-context demos, particularly achieving the best results with 8-shot settings.
6.2 Performance of Hijacking ICL in Classification Task
Our GGI achieves the best hijacking attack performance. While LLMs utilizing ICL show strong performance with clean in-context demos, Table 1 reveal that hijacking attacks significantly undermine their effectiveness. Although baseline methods such as Square, Greedy, and TA are able to partially degrade model performance on the earlier LLMs (e.g., OPT-6.7 and LLaMA-13b), they generally fail to effectively hijack SOTA LLMs, such as LLaMA3.1-8b. Moreover, these methods become inefficient as the number of in-context demos decreases, reflecting their limited efficacy. In contrast, our proposed GGI attack consistently hijacks LLMs to generate the targeted positive sentiment by leveraging just a few shots of adversarially perturbed demos. This is reflected in the near-perfect positive accuracies (approaching 100%) and the drastic collapse of negative accuracies to near 0% across most of the settings, as shown in Table 1. Notably, our GGI remains highly effective even on the SOTA LLMs, demonstrating its superior transferability and robustness compared to baseline attacks.